OpenStudy (anonymous):

solve the following equations: a. 2x^2-4x=6 b. 3^|x-1|=27

OpenStudy (anonymous):

\[a) x^2-2x-3=0\\(x+1)(x-3)=0\\x=-1\text{ or }x=3\]

OpenStudy (anonymous):

b. 3^3=27 so |x-1|=3, x-1=+/- (3) x-1=3 > x=4 x-1=-3 > x=-2 so x=4 or x=-2